OpenStudy (anonymous):

Let y = 25,000(0.94)x represent the purchasing power of $25,000 with an inflation rate of 6%, where x stands for the number of years. The table represents purchasing power in the first four years of inflation, to the nearest dollar. Which is incorrect?

OpenStudy (anonymous):

\[25,000\times (.94)^x\] you job is to compute \[25,000\times (.94)^1 \\25,000\times (.94)^2\\ 25,000\times (.94)^3 \\ 25,000\times (.94)^4\]